Q.1.2 The time taken to complete a web-based questionnaire is normally distributed, with an average time (μ) of 9 minutes and a standard deviation (σ) of 1.55 minutes. What is the probability that a randomly selected person will take: Q.1.2.1 Between 8 and 8.5 minutes to complete the questionnaire? Interpret your answer. (5) Q.1.2.2 Between 8.75 and 9.75 minutes to complete the questionnaire? Interpret your answer.


1
Expert's answer
2021-10-25T13:38:30-0400

"\mu=9 \\

\sigma = 1.55"

1.

"P(8<X<8.5) = P(X<8.5) -P(X<8) \\

= P(Z< \frac{8.5-9}{1.55}) -P(Z< \frac{8-9}{1.55}) \\

= P(Z< -0.322) -P(Z< -0.645) \\

= 0.3737 -0.2594 \\

= 0.1143"

The probability that a randomly selected person will take between 8 and 8.5 minutes to complete the questionnaire is 0.1143.

2.

"P(8.75<X<9.75) = P(X<9.75) -P(Z< 8.75) \\

= P(Z< \frac{9.75 -9}{1.55}) -P(Z< \frac{8.75 -9}{1.55} ) \\

=P(Z < 0.4838) -P(Z< -0.161) \\

= 0.6858 -0.4360 \\

= 0.2498"

The probability that a randomly selected person will take between 8.75 and 9.75 minutes to complete the questionnaire is 0.2498.